# Numerical linked cluster expansions for two-dimensional spin models with continuous disorder distributions

### v1

@@ VIII Summary and discussion
We have shown that NLCEs based on sufficiently large building blocks allow one to obtain accurate low-temperature results for the thermodynamic properties of spin models with continuous disorder distributions in the square lattice. We used three NLCE schemes here, the restricted L, the square, and the rectangle expansions, and carried out the disorder averages directly on the NLCE clusters before computing their weights. We contrasted our results against those obtained using the site expansion, for which it was not possible to control the statistical errors because of the large number of clusters involved in low orders of the expansion. We advance that a similar approach can be used to study models with continuous disorder distributions in other lattice geometries, such as the triangular and kagome lattices, for which a triangle-based expansion is readily available rigol_bryant_07a ; Abdelshafy .
We also showed that for the Ising model with an arbitrary disorder distribution in 1D, NLCEs provide an alternative way (to the traditional transfer matrix method) to obtain the exact analytical result for thermodynamic properties. For the Heisenberg model with a uniform disorder distribution in 1D, we provided evidence that NLCEs allow one to obtain the energy all the way down to the ground state value, and the entropy and specific heat at temperatures that are about two orders of magnitude smaller than the value of $J$ used to set the width of the disorder distribution.
@@ Acknowledgements.
We acknowledge the support of the National Science Foundation Grants No. PHY-2012145 and No. PHY-2012145. The computations were done in the Institute for Computational and Data Sciences’ Roar supercomputer at Penn State.
@@ Appendix A Continuous disorder distributions
In this Appendix, we report additional results obtained using the L, the square, and the rectangle expansions for continuous disorder distributions that have a nonzero mean. We select those distributions, and their corresponding parameters, to be those for which site expansion and Monte Carlo results were reported in Ref. Mulanix_mltidis . The site expansion results in Ref. Mulanix_mltidis were obtained using multi-modal disorder distributions. The results in this appendix allow one to contrast that approach to ours, with which we obtain results at lower temperatures.
@@ A.1 Ising model
For the Ising model, which being a classical model can be studied using Monte Carlo simulations, we consider the bond strengths in Eq. (1) to be of the form $J_{\bf ij}=1+J\,R_{\bf ij}$, with $J=1.5$ and $R_{\bf ij}$ drawn from the uniform distribution $[-1,1]$. In contrast to the case considered in the main text, this distribution exhibits more antiferromagnetic bonds than ferromagnetic ones. In Fig. 10, we show the energy per site vs the temperature from the highest two orders we computed of the L, the square, and the rectangle expansion. We contrast our results to those of Monte Carlo simulations from Ref. Mulanix_mltidis . In agreement with the multi-modal NLCE results for the site expansion reported there (see Fig. 2 in Ref. Mulanix_mltidis ), the results of our direct sums for the L, the square, and the rectangle expansions agree with the Monte Carlo ones at intermediate and high temperature ($T\gtrsim 0.5$). An advantage of the L, the square, and the rectangle expansions over the site expansion results in Ref. Mulanix_mltidis is that the former exhibit direct sums that allow us to compute the energy all the way down to the ground state energy.
@@ A.2 Heisenberg model
Figure 11 shows that the results of the direct sums for the highest order of the L, the square, and the rectangle expansion agree with each other and with the QMC results down to $T\approx 0.5$, which is about one half of the temperature at which the multi-modal NLCE results for the site expansion agree with the QMC ones in Fig. 5 in Ref. Mulanix_mltidis . Like resummations in Ref. Mulanix_mltidis , in the inset in Fig. 11 one can see that Wynn’s resummations of the rectangle expansion extend the agreement of the NLCE results with the QMC ones to lower temperatures, $T\gtrsim 0.3$ in our case.
Figure 11: Energy $E$ per site vs $T$ for the square lattice Heisenberg model with a uniform disorder distribution with nonzero mean (see text). We report results for the highest two orders of the L (L), square ($\square$), and rectangle (R) expansions along with the Monte Carlo (QMC) results reported in Ref. Mulanix_mltidis . (Inset) Results from Wynn’s resummation of the rectangle expansion, Euler’s resummations of the L and square expansions, and the Monte Carlo (QMC) results (same results and legends as in the main panel) reported in Ref. Mulanix_mltidis .

### v2

@@ VIII Summary and discussion
We have shown that NLCEs based on sufficiently large building blocks allow one to obtain accurate low-temperature results for the thermodynamic properties of spin models with continuous disorder distributions in the square lattice. We used three NLCE schemes here, the restricted L, the square, and the rectangle expansions, and carried out the disorder averages directly on the NLCE clusters before computing their weights. We contrasted our results against those obtained using the site expansion, for which it was not possible to control the statistical errors because of the large number of clusters involved in low orders of the expansion. We advance that a similar approach can be used to study models with continuous disorder distributions in other lattice geometries, such as the triangular and kagome lattices, for which triangle-based expansions are readily available rigol_bryant_07a ; Abdelshafy .
We also showed that for the Ising model with an arbitrary disorder distribution in 1D, the linked cluster theorem provides an alternative way (to the traditional transfer matrix method) to obtain the exact analytical result for thermodynamic properties. For the Heisenberg model with a uniform disorder distribution in 1D, we provided evidence that NLCEs allow one to obtain the energy all the way down to the ground-state value, and the entropy and specific heat at temperatures that are about two orders of magnitude smaller than the value of $J$ used to set the width of the disorder distribution.
@@ Acknowledgements.
We acknowledge the support of the National Science Foundation Grants No. PHY-2012145 and No. PHY-2309146. The computations were done in the Institute for Computational and Data Sciences’ Roar supercomputer at Penn State.
@@ Appendix A Continuous disorder distributions
In this Appendix, we report additional results obtained using the L, the square, and the rectangle expansions for continuous disorder distributions that have a nonzero mean. We select those distributions, and their corresponding parameters, to be those for which site expansion and Monte Carlo results were reported in Ref. Mulanix_mltidis . The site expansion results in Ref. Mulanix_mltidis were obtained using multimodal disorder distributions. The results in this appendix allow one to contrast that approach to ours, with which we obtain results at lower temperatures.
@@ A.1 Ising model
For the Ising model, which being a classical model can be studied using Monte Carlo simulations, we consider the bond strengths in Eq. (1) to be of the form $J_{\bf ij}=1+J\,R_{\bf ij}$, with $J=1.5$ and $R_{\bf ij}$ drawn from the uniform distribution $[-1,1]$. In contrast to the case considered in the main text, this distribution exhibits more antiferromagnetic bonds than ferromagnetic ones. In Fig. 10, we show the energy per site vs the temperature from the highest two orders we computed of the L, the square, and the rectangle expansion. We contrast our results to those of Monte Carlo simulations from Ref. Mulanix_mltidis . In agreement with the multimodal NLCE results for the site expansion reported there (see Fig. 2 in Ref. Mulanix_mltidis ), the results of our direct sums for the L, the square, and the rectangle expansions agree with the Monte Carlo ones at intermediate and high temperature ($T\gtrsim 0.5$). An advantage of the L, the square, and the rectangle expansions over the site expansion results in Ref. Mulanix_mltidis is that the former exhibit direct sums that allow us to compute the energy all the way down to the ground-state energy.
@@ A.2 Heisenberg model
Figure 11: Energy $E$ per site vs $T$ for the square lattice Heisenberg model with a uniform disorder distribution with nonzero mean (see text). We report results for the highest two orders of the L (L), square ($\square$), and rectangle (R) expansions along with the Monte Carlo (QMC) results reported in Ref. Mulanix_mltidis . Inset: Results from Wynn’s resummation of the rectangle expansion, Euler’s resummations of the L and square expansions, and the Monte Carlo (QMC) results (same results and legends as in the main panel) reported in Ref. Mulanix_mltidis .
Figure 11 shows that the results of the direct sums for the highest order of the L, the square, and the rectangle expansion agree with each other and with the QMC results down to $T\approx 0.5$, which is about one half of the temperature at which the multimodal NLCE results for the site expansion agree with the QMC ones in Fig. 5 in Ref. Mulanix_mltidis . Like resummations in Ref. Mulanix_mltidis , in the inset in Fig. 11 one can see that Wynn’s resummations of the rectangle expansion extend the agreement of the NLCE results with the QMC ones to lower temperatures, $T\gtrsim 0.3$ in our case.